40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Apr 26 20:43:51 2013, Jenne, Update, LSC, Remeasuring the Schnupp asymmetry 
    Reply  Fri May 3 17:32:59 2013, Jenne, Update, LSC, Remeasuring the Schnupp asymmetry Xarm_SchnuppData_April2013.pngYarm_SchnuppData_April2013.png
Message ID: 8498     Entry time: Fri Apr 26 20:43:51 2013     Reply to this: 8528
Author: Jenne 
Type: Update 
Category: LSC 
Subject: Remeasuring the Schnupp asymmetry 

[Jenne, Annalisa, with guidance from Koji]

We took data to remeasure the Schnupp asymmetry, using the Valera method that Jamie described in elog 4821

1  First, we locked the arms each with their PO(X,Y) signals, to get the alignment of each arm. 

2.  Then, we locked the Xarm with AS55I (Yarm optics, and PRM very misaligned, more than the misalign script).  Since AS55 was saturating, I changed the analog gain from 24dB to 21dB. (After work was completed, the analog gain was put back to the nominal 24dB for both I&Q.)

3.  We set up the Lockin similar to Jamie's description, with a few differences.  We used the same f = 103.1313, but used ampl=10cts.  Sin and cos gain were each 100.  We changed the lowpass filter from 0.1Hz to 0.05Hz (so each measurement had a settling time of at least 20sec).  We were using LSC-Lockin4, so the Lockin matrix was set so Lockin4 was reading from AS55Q, and the LSC output matrix was such that we were actuating on the ETM (X, then Y when we switched arms later).

4.  By hand, we roughly found the zero crossing of the lockin-q output (which corresponded also to zero of the lockin-I, since this is the place where all of the PDH signal was in AS55I, and the lockin was reading AS55Q). 

5.  We took points separated by 0.2 degrees, plus and minus 1 degree from the zero-crossing phase we had found (i.e., for the Xarm, we roughly found the zero crossing at -14.39 deg, so took data from -15.39 to -13.39degrees).  For each phase, we took 5 measurements (using ezcaread), at least 20 seconds apart.  After moving the phase, we waited at least ~40 seconds (watching the lockin outputs on striptool, they had completely settled after 30 or 40 seconds).

6.  We then repeated steps 2, 4 and 5 for the Y arm.  The lockin setup didn't change, except that now we actuate on ETMY.

We did a quick estimate calculation, from our rough zero-crossings to get a rough measurement of the Schnupp asymmetry.  DeltaPhi = (-14.39 -   -19.79) = 5.40 . This gives us (using F_sideband = 5*11066134, the current 11MHz marconi freq) a rough Schnupp asymmetry of 4 cm. 

Analysis to follow.

EDIT, JCD:  The Xarm gain at this time was -0.160, and the Yarm gain was -0.170

ELOG V3.1.3-